Larry is twice as old as his nephew. 8 years ago, he was 3 times as old as his nephew was then. Find their present ages.
Answer by swincher4391(1107)

L = 2N
L -8 = 3*(N -8)
substitute L = 2N
2N - 8 = 3N - 24
-N = -16
N = 16
L = 2*16 = 32